# A Spatiotemporal, Quasi-experimental Causal Inference Approach to Characterize the Effects of Global Plastic Waste Export and Burning on Air Quality Using Remotely Sensed Data

### v2

@@ C.4 Simulation Results and Discussion
First, we observe that the magnitude of the bias ranges between 0.1 and 0.5, while the true ADT ranges between 0.1 and 2.6, with an average value of 1.9. From the MAE plot, we see that the estimator is less stable on the tails (resulting in larger errors), which is to be expected for methods like LLKR. Our intuition is that the asymmetry of the true ADT curve contributes to the bias being higher at higher values of port proximity.
Note that in addition to the finite sample bias of LLKR, which results in imprecise centering of the ADT’s confidence intervals (Hettinger et al., 2025), here we also do not know the correct specification of the propensity score models $\pi_{B}$ and $\pi_{D}$, instead estimating them from the data. Foreshadowing our coverage results, the simulations in Hettinger et al. (2025) indicate that misspecification of the propensity score models has a greater impact on coverage than misspecification of the outcome models.
In Figure S15, we see that neither the Spatial nor Non-spatial Weighted Bootstrap achieves 95% coverage at any value of $\delta$ (port proximity). However, the Spatial Weighted Bootstrap has consistently higher coverage than the Non-spatial method. This makes sense, as the Spatial method generates wider confidence intervals (as shown in Figure S16). Mirroring the results of the MAE plot, we see that the confidence interval coverage tends to be highest in the lower-middle range of $\delta$ (0.25-0.5) for both the Spatial and Non-spatial Weighted Bootstrap.
Now, we compare the differences in results across our four simulation scenarios. Note that both Scenario 2 (increasing $L$) and Scenario 3 (increasing $\sigma_{\text{sp}}$) generate greater similarity in the value of outcomes that are spatially proximate, through different mechanisms, relative to Scenario 1. Whereas, Scenario 4 simply increases the random noise in the outcomes.
In Figures S15 and S16, we see that inducing greater spatial signal (Scenarios 2 and 3) increases the width of the Spatial Weighted Bootstrap’s confidence intervals, resulting in higher coverage. By contrast, the Non-spatial approach responds to increasing $L$ by narrowing its confidence intervals (resulting in lower coverage) and responds to increasing $\sigma_{\text{sp}}$ by widening its confidence intervals (resulting in higher coverage). In Scenario 4, the Spatial approach actually narrows its confidence intervals, likely because increasing $\sigma_{\text{ind}}$ makes it harder to detect the spatially structured error, resulting in under-estimation of the spatial correlation.
In Figure S14, the only parameter configuration that stands out is Scenario 3 (increased $\sigma_{\text{sp}}$), which results in higher MAE (lower precision in the point estimate) across values of $\delta$. In Figure S13, none of the configurations have consistently different bias.

### v3

@@ C.4 Simulation Results and Discussion
First, we observe that the magnitude of the bias ranges between -0.15 and 0.68, while the true ADT ranges between 0.1 and 2.6, with an average value of 1.9. For all scenarios except the one with increased spatial range, the bias is largest on the tails. For the increased spatial range scenario, the bias becomes negative at low values of port proximity.
From the MAE plot, we see that the estimator is less stable on the tails (resulting in larger errors), which is to be expected for methods like LLKR. We observe that while changing the magnitude of noise does not impact the MAE, increasing $\sigma_{sp}$ results in higher MAE across all values of port proximity. To a lesser extent, increasing the spatial range results in a smaller MAE.
In Figure S15, we see that the Spatial Weighted Bootstrap has consistently higher coverage than the Non-Spatial Weighted Bootstrap, in part due to having wider confidence intervals as shown in Figure S16. Mirroring the results of the MAE and Bias plots, both approaches have higher than 95% coverage in the middle of the port proximity distribution, despite their confidence intervals narrowing. However, at low values of port proximity both methods have less-than-nominal coverage. At high values of port proximity, the Spatial Weighted Bootstrap outperforms the Non-Spatial approach, due to having wider confidence intervals. These results are stable across different numbers of bootstrap replicates (Figure S17). Comparing across scenarios: we see that adding more spatially-structured noise increases the width of the confidence intervals. By contrast, the confidence intervals narrow slightly when the spatial range is increased. Similar to MAE, increasing the magnitude of random noise does not appear to change the interval widths.
To elucidate the extent to which under-coverage on the tails is driven by the finite sample bias of LLKR, we did as suggested by Hettinger et al. (2025a) and evaluated the coverage after subtracting out the average bias. Figure S18 indicates that especially at higher values of port proximity, the LLKR bias is playing a substantial role in the coverage for both the Spatial and Non-Spatial Weighted Bootstrap approaches. At lower values of port proximity, the coverage gets pulled closer to 95% for all scenarios except that with increased spatial range.
@@
Figure S17: Coverage and width of the confidence intervals when using 100 vs 200 bootstrap replicates.
Figure S18: Coverage of the 95% confidence intervals in our simulations, after removing the average bias. The dashed line indicates the desired 95% coverage.